Tickets for adults are 5.50 and tickets for children are 3.50. How many of each kind of ticket was purchased if 21 tickets were bought for 83.50?
Answer by checkley77(12569)

5.5X+3.50(21-X)=83.50
5.5X+73.5-3.50X=83.50
2X=83.50-73.50
2X=10
X=10/2
X=5 ADULT TICKETS WERE SOLD.
21-5=16 CHILDREN TICKETS WERE SOLD.
PROOF:
5.50*5+3.50*16=83.50
27.50+56.00=83.50
83.50=83.50